Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A miniature roulette wheel is divided into 10 equal sectors,

Expert replies
by AAPL » Thu Jan 24, 2019 5:17 am
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

Veritas Prep

A miniature roulette wheel is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly determines the winning sector by settling in that sector. If the wheel is spun three times, approximately what is the probability that the product of the three winning sectors’ integers will be even?

A. 88%
B. 75%
C. 67%
D. 63%
E. 50%

OA A
Join the discussion
Source: — Problem Solving |

AAPL wrote:Veritas Prep

A miniature roulette wheel is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly determines the winning sector by settling in that sector. If the wheel is spun three times, approximately what is the probability that the product of the three winning sectors’ integers will be even?

A. 88%
B. 75%
C. 67%
D. 63%
E. 50%
\[?\,\,\, \cong \,\,\,1 - P\left( {\underbrace {{\text{all}}\,\,{\text{odd}}\,\,{\text{sectors}}}_{{\text{unfavorable}}}} \right)\]
$$P\left( {{\rm{all}}\,\,{\rm{odd}}\,\,{\rm{sectors}}} \right)\,\,\,\mathop = \limits^{{\rm{independency}}} \,\,\,{1 \over 2} \cdot {1 \over 2} \cdot {1 \over 2} = {1 \over 8}$$
\[? = 1 - \frac{1}{8} = \frac{7}{8} = \underleftrightarrow {7 \cdot \frac{{0.25}}{2} > 7 \cdot \frac{{0.24}}{2}} = 84\% \]

(Please note that our approximation is good enough for the alternative choices offered.)


We follow the notations and rationale taught in the GMATH method.

Regards,
Fabio.
Fabio Skilnik :: GMATH method creator ( Math for the GMAT)
English-speakers :: https://www.gmath.net
Portuguese-speakers :: https://www.gmath.com.br
Join the discussion

by [email protected] » Thu Jan 24, 2019 10:10 am
Hi All,

We're told that a miniature roulette wheel is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly determines the winning sector by settling in that sector and the wheel is spun three times. We're asked for the approximate probability that the PRODUCT of the three winning sectors’ integers will be even. This question is based on a few Number Properties and can be solved in a couple of different ways, depending on how you want to organize your work.

To start, the following Number Properties are worth knowing:
(Even)(Even) = Even
(Even)(Odd) = Even
(Odd)(Odd) = Odd

When multiplying integers, to end up with a PRODUCT that is EVEN, we need AT LEAST ONE of the integers to be Even. If ALL of the integers are ODD, then the product will be ODD. On this wheel, the probability of landing on an even number or landing on an odd number is the same: 1/2. We can either calculate the probability of ending up with an even product or the probability of ending with an odd product (and then subtracting that fraction from 1). As it stands, there are only 8 possible arrangements of the 3 winning results:

(Even)(Even)(Even) = Even
(Even)(Even)(Odd) = Even
(Even)(Odd)(Even) = Even
(Odd)(Even)(Even) = Even
(Even)(Odd)(Odd) = Even
(Odd)(Even)(Odd) = Even
(Odd)(Odd)(Even) = Even
(Odd)(Odd)(Odd) = Odd

The probability of ending with an EVEN product is 7/8 = .875 = 87.5% --> approximately 88%

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Scott@TargetTestPrep » Thu Jan 31, 2019 5:39 pm
AAPL wrote:Veritas Prep

A miniature roulette wheel is divided into 10 equal sectors, each bearing a distinct integer from 1 to 10, inclusive. Each time the wheel is spun, a ball randomly determines the winning sector by settling in that sector. If the wheel is spun three times, approximately what is the probability that the product of the three winning sectors’ integers will be even?

A. 88%
B. 75%
C. 67%
D. 63%
E. 50%

OA A
In order to have an even product, at least one of the numbers must be even.

If the 1st number is even, then it doesn’t matter what the 2nd and 3rd numbers are. So P(1st is even) = 1/2 x 1 x 1 = 1/2.

If the 1st number is odd and the second number is even, then it doesn’t matter what the 3rd numbers is. So P(1st is even and 2nd iis odd) = 1/2 x 1/2 x 1 = 1/4.

If the 1st number is odd and the second number is odd, then the 3rd numbers must be even so that the product is even. So P(1st is even and 2nd is odd and 3rd is odd) = 1/2 x 1/2 x 1/2 = 1/8.

Since all three cases above are mutually exclusive, the probability that the product of the three winning sectors’ integers will be even is 1/2 + 1/4 + 1/8 = 7/8 = 87.5% or approximately 88%.

Alternate solution:

In order to have an even product, at least one of the numbers must be even. The only time we don’t have any even number is if the ball lands on an odd number for all three spins. In other words, the ball lands on 3 odd numbers in 3 spins. The probability that this will occur is 1/2 x 1/2 x 1/2 = 1/8. So the probability that this will not occur, i.e., on at least one spin the ball will land on an even number, is 1 - 1/8 = 7/8 = 87.5%, or approximately 88%.

Answer: A

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion